OpenStudy (anonymous):

HELP!! please 8th grade i will fan and medal. Events A and B are mutually exclusive, with P(A) = 0.65 and P(B) = 0.14. What is P(A or B)? A. 0.091 B. 0.215 C. 0.51 D. 0.79

OpenStudy (anonymous):

@Yothatboy

OpenStudy (anonymous):

p(A or B)=p(A)+p(b)=0.14+0.65=0.79

OpenStudy (anonymous):

OH! Thank you so much!!